# The DAMIC-M Low Background Chamber

### v1

@@ 5.4 Background measurements
Several runs dedicated to the study of the radioactive background were performed in 2022 and 2023. These runs had various configurations, typically no pixel binning [^6] Pixel binning is an operating mode of the CCD where the charge of several pixels is summed before being read out. The charge can be summed only in the vertical or horizontal direction, or in both directions. was used for the acquired images and the whole CCD was read out after long exposures (up to half a day).
@@
An analysis of $\alpha$ rates and coincidences, e.g. $\beta-\beta$ for 32Si and 210Pb, on the surface and in the bulk of the 6k×$\times$1.5k CCDs is under way. Our studies follow previous work done for DAMIC at SNOLAB [67]. The measurement of contaminants inside the silicon bulk is of primary importance for the final DAMIC-M detector, because the CCDs in DAMIC-M will be made from the same silicon ingot as 6k×$\times$1.5k CCDs installed in the LBC.
@@ 6.1 Pressure and Temperature
Unlike pressure gauges, temperature sensors are monitored continuously. The RTD in the cold mass table, see Fig. 3, is kept at a constant temperature by using a 100 W heater located 2.5 cm next to it. The reading of an RTD epoxied in a cold finger, where it is well shielded by lead plates from CCDs, is about 2 K higher. An RTD was kept temporarily on the CCD box during the detector commissioning and measured 4 K higher than the cold finger. This RTD was then removed to suppress the radioactive background. The highest temperature gradient – about 30 K – is between the cold head and the cold mass table and it is due to a polyimide tape, which provides an electrical isolation but reduces thermal contact. The temperature of the CCDs has been measured through independent thermal tests and is typically $\sim$5 K higher than the temperature measured by an RTD epoxied on the outside of the CCD box.
Over 63 consecutive days of data taking (described in Refs. [24, 31]), the cold mass table temperature increased from 116.9 K to 117.2 K, see Fig. 7. This increase may be explained either by heating the cryostat by a flow of air warmed by the cryocooler body when the external shielding is closed, or relaxing the connection between cold copper parts. As is shown in Refs. [31], the dark current did not increase during the same period.
@@ 6.2 CCD Performance
The LBC serves as a test setup for the DAMIC-M project and allows us to perform various studies of operating parameters for different CCD formats and electronics. An extensive optimization program for clock and bias voltages, and timings were carried out. These tests were performed with originally installed 6k×$\times$4k CCDs and then the CCD modules with 6k×$\times$1.5k devices. The latter is of particular importance, because very similar CCD modules will be installed in the full-scale DAMIC-M detector.
@@
The first studied noise levels were $\sigma_{1}\sim 10$ e- which are higher than in DAMIC at SNOLAB and test setups located in different institutions. After implementing noise mitigation steps, the noise level decreased to an acceptable level of $\sigma_{1}\sim 7$ e- in images taken with ARC CCD controllers in Setup 1. The averaging in the skipper amplifier reduces any remaining high frequency noise contributions.
@@
We performed additional tests to characterize installed skipper CCDs. These include study the temperature dependence of dark current and the increase in dark current for different levels of ionizing radiation from the external 60Co source.
@@ 8 Acknowledgements
We are grateful to LSM and its staff for support through underground space, logistical and technical services.

### v2

@@ 5.4 Background measurements
Several runs dedicated to the study of the radioactive background were performed in 2022 and 2023. To measure the background, pixel binning [^6] Pixel binning is an operating mode of the CCD where the charge of several pixels is summed before being read out. The charge can be summed only in the vertical or horizontal direction, or in both directions. was not used for the acquired images. Instead, the entire CCD was read out after a long exposure (up to half a day). This approach was chosen to be able to effectively utilize the coincidence method for measuring bulk radioactive contamination, benefiting from the excellent spatial resolution of the CCDs.
@@
An analysis of $\alpha$ decay rates and spatially coincident events, such as $\beta-\beta$ for 32Si and 210Pb, is currently underway for the surface and bulk of the 6k×$\times$1.5k CCDs. Our studies follow previous work done for DAMIC at SNOLAB [69]. Measuring contaminants within the silicon bulk is crucial for the final DAMIC-M detector, as the CCDs used in DAMIC-M will be made from the same silicon ingot as those currently installed in the LBC.
@@ 6.1 Pressure and Temperature
Unlike pressure gauges, temperature sensors are monitored continuously. The RTD in the cold mass table, see Fig. 3, is kept at a constant temperature by using a 100 W heater, see Sec. 4.2. The reading of an RTD epoxied in a cold finger, where it is well shielded by lead plates from CCDs, is about 2 K higher. An RTD was kept temporarily on the CCD box during the detector commissioning and measured 4 K higher than the cold finger. This RTD was then removed to suppress the radioactive background. The highest temperature gradient – about 30 K – is between the cold head and the cold mass table and it is due to a polyimide tape, which provides an electrical isolation but reduces thermal contact. The temperature of the CCDs has been measured through independent thermal tests and is typically $\sim$5 K higher than the temperature measured by an RTD epoxied on the outside of the CCD box.
Over 63 consecutive days of data taking (described in Refs. [24, 31]), the cold mass table temperature increased only slightly from 116.9 K to 117.2 K, see Fig. 7. This increase may be explained either by heating the cryostat by a flow of air warmed by the cryocooler body when the external shielding is closed, or relaxing the connection between cold copper parts. This slow temperature increase was no issue during the data taking, because it did not cause any change of measured dark current as is shown in Refs. [31].
@@ 6.2 CCD Performance
The LBC serves as a test setup for the DAMIC-M project and allows us to perform various studies of operating parameters for different CCD formats and electronics. An extensive optimization program for clock and bias voltages, and timings were carried out. These tests were performed with the originally installed 6k×$\times$4k CCDs and then the CCD modules with 6k×$\times$1.5k devices. The latter is of particular importance, because very similar CCD modules will be installed in the full-scale DAMIC-M detector.
@@
The first studied noise levels were $\sigma_{1}\sim 10$ e- which are higher than in DAMIC at SNOLAB and test setups located in different institutions. After implementing noise mitigation steps (e.g. improved electrical grounding and better timing of voltage clocks), the noise level decreased to an acceptable level of $\sigma_{1}\sim 7$ e- in images taken with ARC CCD controllers in Setup 1. The averaging in the skipper amplifier reduces any remaining high frequency noise contributions.
@@
We performed additional tests to characterize installed skipper CCDs. These include study the temperature dependence of dark current and the increase in dark current for different levels of ionizing radiation from the external 60Co source. The analysis is ongoing and will be published elsewhere.
@@ 8 Acknowledgements
We would like to thank the Modane Underground Laboratory and its staff for support through underground space, logistical and technical services. LSM operations are supported by the CNRS, with underground access facilitated by the Societe Francaise du Tunnel Routier du Frejus